OpenStudy (anonymous):

Chapter 5:Indices and Logarithms Double checking Simplify the following @ganeshie8

OpenStudy (anonymous):

\[2^{n}-2^{n+1}\]

OpenStudy (anonymous):

\[(2^n-2^n~\times~2^{-1})\]

OpenStudy (anonymous):

\[2^n(1-1-1)\]

OpenStudy (anonymous):

\[-2^n\]

ganeshie8 (ganeshie8):

\[2^n- 2^{n+1} = 2^n - 2\cdot 2^n = 2^n(1-2) = -2^n\]

OpenStudy (anonymous):

Thnx @ganeshie8

ganeshie8 (ganeshie8):

yw!